is x= -5/2 a solution of the inequality 5x-4<_3 (x-7)?
Answer by stanbon(75887)

is x= -5/2 a solution of the inequality 5x-4<_3 (x-7)?
---
Substitute to see if you get a true statement.
----
5(-5/2) -4 <= 3(-5/2-7)
----
-25/2 -4 <= 3(-19/2)
-----
-33/2 <= -57/2
---
False
---
So x=-5/2 is not a solution for the inequality.
==================================================
Cheers,
Stan H.